The hour hand of a huge clock is 100cm long. What is the

circumferential speed of the tip of this hand in cm/hr.?


1
Expert's answer
2021-09-21T12:28:35-0400

Length of hour hand

L=10cm

Distance covered by hour hand in 1hour

"D=2\pi L=2\pi ×100\\=200\pi cm"

circumferential speed of the tip

"V=\frac{D}{t}=\frac{200\pi}{1}cm/ hr\\=200\pi cm/hr\\=628.32 cm/hr"